# Testing the Origin of Hot Jupiters with Ariel

### v1

@@ Abstract
In spite of their long detection history, the origin of hot Jupiters remains to be resolved. While multiple dynamical evidence suggests high-eccentricity migration is most likely, conflicts remain when considering hot Jupiters as a population in the context of warm and cold Jupiters. Here, we turn to atmospheric signatures as an alternative mean to test the origin theory of hot Jupiters, focusing on population level trends that arise from post-formation pollution, motivated by the upcoming Ariel space mission whose goal is to deliver a uniform sample of exoplanet atmospheric constraints. We experiment with post-formation pollution by planetesimal accretion, pebble accretion, and disk-induced migration and find that an observable signature of post-formation pollution is only possible under pebble accretion in metal-heavy disks. If most hot Jupiters arrive at their present orbit by high-eccentricity migration while warm Jupiters emerge largely in situ, we expect the atmospheric water abundance of hot Jupiters to be significantly elevated compared to warm Jupiters, at a level that is easily detectable by Ariel transit survey. We provide suggestions for future comparative atmospheric characterization between hot Jupiters and wide-orbit directly imaged planets to elucidate the properties of the dust substructures in protoplanetary disks.
@@ 1 Introduction
Hot Jupiters are one of the first kind of exoplanet ever detected. Yet, their origin remains debated. Broadly, three origin hypotheses have been proposed in the literature, including in situ formation, disk-induced migration, and high-eccentricity migration, all of which have their successes and shortcomings (see R. I. Dawson & J. A. Johnson, 2018, for a review). Among these three, in situ formation is the least likely given the need to coagulate cores of mass (e.g., R. R. Rafikov, 2006; E. J. Lee et al., 2014; A.-M. a. Piso et al., 2015; V. Savignac & E. J. Lee, 2024) which is prohibitively large at such short orbital distances (e.g., B. Bitsch et al., 2018; J. Fung & E. J. Lee, 2018).
@@
An alternative path to resolving the origin of hot Jupiters is to use chemical abundances. Early and recent studies have focused on the expected atmospheric composition of hot Jupiters depending on their initial formation locations (e.g., N. Madhusudhan et al., 2014; A. D. Schneider & B. Bitsch, 2021; A. B. T. Penzlin et al., 2024), based on the radial variation in chemical content in protoplanetary disks (e.g., K. I. Öberg et al., 2011; P. Mollière et al., 2022). These studies generally expect high C/O ratio as a signature of formation beyond the iceline as the water molecules are expected to be in solid form and gas giants, by definition, are mostly gas. Pollution of the envelope by continued solid accretion after formation can complicate this picture, however, as the infalling solids are expected to evaporate within the envelope (e.g., E. Vlahos et al., 2024). Support for post-formation pollution of gas giants may be found in the high atmospheric metallicities of our own Jupiter (e.g., K. I. Öberg & R. Wordsworth, 2019) and HR 8799 planets (E. Nasedkin et al. 2024; Ruffio et al. submitted), as well as low C/O ratio of Pic b (GRAVITY Collaboration et al., 2020).
While the collection of dynamical evidence suggests distinct origins of hot vs. warm Jupiters, a clear alternative test from atmospheric studies remains to be proposed. In this paper, we aim to develop population-level atmospheric signatures (i.e., trends) as a probe of disk-induced vs. high-eccentricity migration for hot and warm Jupiters. We are particularly motivated by the upcoming European Space Agency’s Ariel mission that promises to deliver a uniform sample of gas giants over a range of parameter space, ideally suited to identifying trends rather than a deep dive into individual planets.
@@ 2 Theories
We first construct theoretically expected trends in atmospheric metallicity, focusing on the deviation of the planetary metallicity from the stellar value due to post-formation pollution. Our assumptions are: 1) the giants inherit the stellar metallicity at formation; and 2) the infalling material is well-mixed with the envelope by convection. Following these assumptions, we consider observable atmospheric metallicity as the mass of solids accreted divided by the mass of the planet, fixed to one Jupiter mass for simplicity.111We have implicitly assumed the mass of the initial rocky core that triggered the nucleation of a giant is significantly smaller than one Jupiter mass, which is valid by the definition of a gas giant.
@@ 2.1 Pollution by Solid Accretion
We consider three different paths of pollution: in situ planetesimal accretion, in situ pebble accretion, and migration. In reality, the modes of accretion are likely mixed, but we keep the three paths distinct for clarity. In all cases, we consider a giant planet of mass and radius to have already formed and be simply further accreting the remaining solids in the disk.
@@
where is the local solid disk surface density, is the Keplerian orbital frequency, is the gravitational constant, is the mass of the star, is the orbital distance, and is the radius of the planet. We do not consider gravitational focusing as it is unclear a priori what the typical random velocity of the planetesimals should be. We therefore consider the simplest case of geometric cross section.
@@
As the orbiting planet tidally perturbs the gas, the gas torques back on the planet. Any non-zero net torque will cause the planet to migrate inwards or outward, depending on the sign of the net torque. The exact behavior (both the direction and the magnitude) of the migration depends very sensitively on the disk structure, including its turbulent and radiative properties (e.g., S. Paardekooper et al., 2023). Here, we adopt Type I migration modified for gap opening presented by K. D. Kanagawa et al. (2018) which likely places an upper limit on the pollution rate by migration. The torque is related to the orbital angular momentum of the planet via
@@
with , the Shakura-Sunyaev turbulent alpha.222While the is constrained to be low in the outer regions of the disk (; e.g., C. Pinte et al. 2015, K. M. Flaherty et al. 2017, E. J. Lee 2024), it is unclear whether would stay as low at short orbital periods that are relevant to Ariel. In writing equation 7, we ignored order unity numerical coefficient (2.5 for our choice of and ). We discuss below in Section 3 how accounting for this coefficient would not change our final result. Planets were initially placed between 0.5 and 50 AU and we evolve their inward migration by integrating equation 6.
@@
where is set by equation 2, is a depletion factor we vary between 0.001 and 1, and Myr is the disk lifetime (e.g., E. E. Mamajek, 2009).333While the more improved disk lifetime is closer to 7–8 Myr (e.g., A. Michel et al., 2021), Jupiters are more likely to appear around more massive stars which tend to have shorter-lived disks (e.g., Á. Ribas et al., 2015) so we use 2 Myr as our nominal value.
@@ 2.2.1 Dust Replenishment
Our model setup implicitly treats the planet feeding zone to be isolated (i.e., no replenishment of solids). For planetesimal accretion, unless there are neighboring planets that can systematically scatter the planetesimals in, there is no a priori reason why a given feeding zone would be replenished. For pebble accretion, the picture is more complicated as the pebbles drift in aerodynamically and the drift timescale is shorter than pebble accretion time by factors of up to 10 (e.g., J. W. Lin et al., 2018; Y. Chachan & E. J. Lee, 2023). One way to stop this radial drift-in is to establish a pressure trap (e.g., P. Pinilla et al., 2012). While the exact source of the pressure trap remains debated—the most popular being perturbation by an embedded planet—it is a natural mechanism to explain both the longevity of dust disks and the observed concentric ring-like structures in protoplanetary disks (see, e.g., J. Bae et al., 2023, for a review).
@@
Given the range of possibilities, we explore the two limiting cases. Our isolated in situ pebble accretion corresponds to the case of a perfect trap created by the accreting Jupiter. We introduce another limiting case of a trap that is so inefficient that the trap is functionally non-existent. In this latter case, the planet has access to all the initial dust material exterior to its orbit:
@@ 2.2.2 Water Abundance
So far, we have computed the total mass of the metals that would be accreted onto a giant planet through post-formation pollution. Creating a large uniform sample of atmospheric measurements is most amenable for volatile species such as water which is more easily detectable than tracers of refractory species such as sulfur or nitrogen. Treatment of water abundance depends on the location at which the solids were accreted. Inside the iceline, water is in gaseous form, so post-formation accretion of solids would not impact water abundance, hence the water abundance is set by the initial solar value, which we take as 0.48 (oxygen mass fraction among all metals; N. Grevesse & A. Noels 1993) times 0.02 (our adopted solar metallicity), irrespective of pollution parameters. Outside the iceline, water is in solid form, so we estimate the water abundance by multiplying the total accreted metal mass by 0.48, assuming all the oxygen to be incorporated into water for simplicity (see, e.g., K. Heng et al., 2016, their Figure 3, “cold” case). We take the water iceline at 160 K which is 0.86 AU (291 days) based on our adopted disk temperature profile (equation 3). For more direct comparison with observable, we further multiply our mass ratio by 2.3/18 (solar metallicity gas mean molecular weight / mass of water molecule) to obtain number ratio.
@@ 3 Results
The predicted trends of metallicity (accreted solid mass over Jupiter mass) following post-formation pollution by planetesimal accretion, pebble accretion, and disk-induced migration are found in Figures 2, 3, and 4 respectively. The abundances shown in these figures represent what is added to the planet after formation; therefore, they do not represent the absolute atmospheric abundances—in order to obtain the absolute abundances, our results would have to be added to the metallicities the planets inherited at formation from the disk, which we assume to be stellar for simplicity. The implication is that any additive metallicity that falls below the stellar value (fixed to be 0.02 for simplicity) would simply not show up in observations as a signature of pollution. We can therefore rule out all pollution methods except pebble accretion with high and high to be interesting.
@@
which drops with orbital distance. In other words, the key difference is that in planetesimal accretion the planet’s accretion cross section stays fixed, whereas in local pebble accretion the accretion cross section is bigger farther from the star.
Likewise, under disk-induced migration, metallicity declines with the final orbital period. By definition, if the final orbital period is long, there has been minimal radial migration and the planet spends most time where is lower, leading to overall lower metallicity by pollution. Figure 4 shows that the behavior of metallicity with respect to (gas disk depletion factor) is non-monotonic. While the total amount of metals accreted for planets migrating in gas-full () disks is larger, these planets also migrate the fastest and reach the shortest final orbital period, causing an overall leftward shift in the metallicity-final-period diagram (e.g., planets that initially start at 105 days with the lowest metallicity park at 103 days) with the planets that attain highest metallicity all pile up at the inner edge of 10 days.
@@ 4.1 Accuracy and Precision of Tier 2 Transit Spectroscopy Retrievals
To simulate Ariel observations using real targets, we use the July 2024 version of the Ariel Mission Candidate Sample (MCS) presented by B. Edwards & G. Tinetti (2022) and available on GitHub. We use the optimistic version of the MCS, which includes both confirmed and candidate planets. We conservatively adopt the easiest targets observable in a cumulative observing time of 1 year, where we assume that each transit observation takes 3 the transit duration and each target is re-observed until it reaches Tier 2 signal-to-noise. Our mock survey includes 265 Jupiters.
Ariel’s targets are hot and warm Jupiters around FGK and early M type dwarfs so they are all within the iceline. In the simplest baseline scenario of minimal migration after pollution, we expect the water abundance to be flat with respect to orbital period set to the stellar value irrespective of the existence/efficiency of a dust trap, as shown in Figure 5. We initially assign this stellar water abundance to each planet (adding stellar [Fe/H] to the solar value we adopt) and then randomly shift the abundance by 0.100.07 dex. Figure 6 depicts simulated Ariel retrievals of 265 Tier 2 transmission spectroscopy targets after 1 year of cumulative observation time with Ariel. A fit through all the planets (i.e., raw data without binning) recovers the expected flat H2O abundance with respect to orbital period.
@@
Under high-eccentricity migration, a gas giant first completes its formation at large orbital separations (beyond the iceline). After the disk dissipates, the giant is excited to a highly eccentric orbit by neighboring planetary or stellar perturbers, then tidally circularizes to a tight orbit with orbital period less than 10 days. Under all circumstances, giants that complete their pollution beyond the iceline are expected to have a significantly water-enriched envelope as compared to those that remain inside the iceline (because solid pollution adds extra water only beyond the iceline). We therefore expect the signature of high-eccentricity migration to show up as a sudden jump in water abundance ratio for hot Jupiters compared to warm Jupiters—insofar as most warm Jupiters did not undergo similar dynamical upheaval (e.g., M. Rice et al., 2022b; X.-Y. Wang et al., 2024).444For the purposes of the simulation, we assume that the migrating Jovian planets have already reached an orbital period on the order of days. Additionally, we assume that the metallicity of the planet remains constant throughout the high-eccentricity migration.
@@
Our goal is to assess the ability of Ariel survey at the Tier 2 level to discern the aforementioned effect of high-eccentricity migration on the measurable water abundance of hot Jupiters. To simulate the migration of a giant from beyond to inside the iceline, we simply move the planets from beyond 1 AU down to 1–4 days where we observe hot Jupiters. We generate 250 fictitious exoplanets, as this is roughly the number of Tier 2 transmission spectroscopy targets that Ariel could observe within a year. These mock planets have arbitrary initial orbital distances between 0.08 and 10 AU (for stellar masses drawn uniformly between 0.4 and 3 encompassing 97.6% of the Jovian hosts in the MCS), water abundance set by model expectation in Figure 5 corrected to non-solar abundance drawing from the stellar [Fe/H] distribution in the MCS sample, as well as uncertainties drawn from Q. Changeat & K. H. Yip (2023) set to 0.100.07 dex. The 10 AU upper limit to the orbital distance is motivated by the observed peak of gas giant occurrence rate at 1–10 AU (B. J. Fulton et al., 2021).
@@
Figure 7 illustrates the resulting water abundance of hot vs. warm Jupiters. As expected, polluted hot Jupiters that undergo high-eccentricity migration from beyond the iceline are expected to be significantly more water-enriched than warm Jupiters at 0.5–1.0 dex level, which is enough to be detectable by Ariel, and visible by eye when the data are binned. The scenario with no dust traps predicts the hot Jupiters to be noticeably more water-rich than those with perfect dust traps simply because they have access to more solids (see Figure 5). Much of the scatter in true H2O value for hot Jupiters comes from our random distribution of the initial orbital distance without any correlation to the final location in addition to the scatter in stellar metallicity. While a more detailed orbital evolution may beget further trends with e.g., orbital period among the hot Jupiters, we defer such studies to future investigation.

### v3

# Testing the Origin of Hot Jupiters with Atmospheric Surveys

@@ Abstract
In spite of their long detection history, the origin of hot Jupiters remains to be resolved. While multiple dynamical evidence suggests high-eccentricity migration is most likely, conflicts remain when considering hot Jupiters as a population in the context of warm and cold Jupiters. Here, we turn to atmospheric signatures as an alternative mean to test the origin theory of hot Jupiters, focusing on population level trends that arise from post-formation pollution, motivated by the upcoming Ariel space mission whose goal is to deliver a uniform sample of exoplanet atmospheric constraints. We experiment with post-formation pollution by planetesimal accretion, pebble accretion, and disk-induced migration and find that an observable signature of post-formation pollution is only possible under pebble accretion in metal-heavy disks. If most hot Jupiters arrive at their present orbit by high-eccentricity migration while warm Jupiters emerge largely in situ, we expect the atmospheric water abundance of hot Jupiters to be significantly elevated compared to warm Jupiters. We report on the detectability of such signatures and further provide suggestions for future comparative atmospheric characterization between hot Jupiters and wide-orbit directly imaged planets to elucidate the properties of the dust substructures in protoplanetary disks.
@@ 1 Introduction
Hot Jupiters are one of the first kind of exoplanet ever detected. Yet, their origin remains debated. Broadly, three origin hypotheses have been proposed in the literature, including in situ formation, disk-induced migration, and high-eccentricity migration, all of which have their successes and shortcomings (see R. I. Dawson & J. A. Johnson, 2018, for a review). Among these three, in situ formation (e.g., K. Batygin et al., 2016) is the least likely given the need to coagulate cores of mass (e.g., R. R. Rafikov, 2006; E. J. Lee et al., 2014; A.-M. a. Piso et al., 2015; V. Savignac & E. J. Lee, 2024) which is prohibitively large at such short orbital distances by e.g., pebble accretion (e.g., B. Bitsch et al., 2018; J. Fung & E. J. Lee, 2018). While continued planetesimal accretion with enough solids in the background disk could generate 10 cores on timescales 104 yr at orbital periods of a few days (see e.g., E. J. Lee et al., 2014, their Section 1.1), observations find that the vast majority of 10 planets at these short orbital periods stay as sub-Neptunes rather than blowing up into hot Jupiters in situ (e.g., F. Fressin et al., 2013; S. Dong & Z. Zhu, 2013; E. A. Petigura et al., 2018).
@@
An alternative path to resolving the origin of hot Jupiters is to use chemical abundances. Early and recent studies have focused on the expected atmospheric composition of hot Jupiters depending on their initial formation locations (e.g., N. Madhusudhan et al., 2014; A. D. Schneider & B. Bitsch, 2021; A. B. T. Penzlin et al., 2024), based on the radial variation in chemical content in protoplanetary disks (e.g., K. I. Öberg et al., 2011; P. Mollière et al., 2022). These studies generally expect high C/O ratio as a signature of formation beyond the iceline as the water molecules are expected to be in solid form and gas giants, by definition, are mostly gas (i.e., gas-phase C/O is expected to be high so it follows that gas-dominated accretion leads to high C/O). Pollution of the envelope by continued solid accretion after (or during) formation can complicate this picture, however, as the infalling solids are expected to evaporate within the envelope (e.g., O. Mousis et al., 2009; J. J. Fortney et al., 2013; C. Mordasini et al., 2016; S. Shibata et al., 2020; E. Vlahos et al., 2024). Support for post-formation pollution of gas giants may be found in the high atmospheric metallicities of our own Jupiter (e.g., K. I. Öberg & R. Wordsworth, 2019) and HR 8799 planets (E. Nasedkin et al. 2024; J-B. Ruffio et al. submitted), as well as low C/O ratio of Pic b (GRAVITY Collaboration et al., 2020). At short orbital periods, supersolar metallicity is reported for WASP-15b which is a spin-orbit misalignd hot Jupiter (J. Kirk et al., 2025), and interior structure models infer high (superstellar) bulk metallicity of warm Jupiters (D. P. Thorngren et al., 2016).
While the collection of dynamical evidence suggests distinct origins of hot vs. warm Jupiters, a clear alternative test from atmospheric studies remains to be proposed. In this paper, we aim to develop population-level atmospheric signatures (i.e., trends) as a probe of disk-induced vs. high-eccentricity migration for hot and warm Jupiters. We are particularly motivated by the upcoming European Space Agency’s Ariel mission that promises to deliver a uniform sample of gas giants over a range of parameter space, ideally suited to identifying trends rather than a deep dive into individual or a small selection of planets.111While we benchmark all our calculations and discussions to the projected capabilities of Ariel, our results will be broadly applicable for any survey design comparable to Ariel. Our investigation is therefore complementary to ongoing James Webb Space Telescope (JWST) programs such as BOWIE-ALIGN that specifically targets a handful (8) of hot Jupiters around stars above the Kraft break (J. Kirk et al., 2024) and JWST Cycle 4 GO 9025 that specifically targets 5 warm Jupiters that are spin-orbit aligned (P. Gao et al., 2025).
@@ 2 Theories
We first construct theoretically expected trends in atmospheric metallicity, focusing on the deviation of the planetary metallicity from the stellar value due to post-formation pollution. Our assumptions are: 1) the giants inherit the stellar metallicity at formation; and 2) the infalling material is well-mixed with the envelope by convection. Following these assumptions, we consider observable atmospheric metallicity as the mass of solids accreted divided by the mass of the planet, fixed to one Jupiter mass for simplicity.222We have implicitly assumed the mass of the initial rocky core that triggered the nucleation of a giant is significantly smaller than one Jupiter mass, which is valid by the definition of a gas giant.
@@ 2.1 Pollution by Solid Accretion
We consider three different paths of pollution: in situ planetesimal accretion, in situ pebble accretion, and migration. In reality, the modes of accretion are likely mixed, but we keep the three paths distinct for clarity. In all cases, we consider a giant planet of mass and radius to have already formed and be simply further accreting the remaining solids in the disk. Our results can be easily scaled to planets of different masses whose implication we discuss in Section 5.
@@
where is the local solid disk surface density, is the Keplerian orbital frequency, is the gravitational constant, is the mass of the star, is the orbital distance, and is the radius of the planet. We do not consider gravitational focusing as it is unclear a priori what the typical random velocity of the planetesimals should be. We therefore consider the simplest case of geometric cross section, which is a conservative limit.
@@
As the orbiting planet tidally perturbs the gas, the gas torques back on the planet. Any non-zero net torque will cause the planet to migrate inward or outward, depending on the sign of the net torque. The exact behavior (both the direction and the magnitude) of the migration depends very sensitively on the disk structure, including its turbulent and radiative properties (e.g., S. Paardekooper et al., 2023). Here, we adopt Type I migration modified for gap opening presented by K. D. Kanagawa et al. (2018) which likely places an upper limit on the pollution rate by migration. The torque is related to the orbital angular momentum of the planet via
@@
with , the Shakura-Sunyaev turbulent alpha.333While the is constrained to be low in the outer regions of the disk (; e.g., C. Pinte et al. 2015, K. M. Flaherty et al. 2017, E. J. Lee 2024), it is unclear whether would stay as low at short orbital periods that are relevant to Ariel. In writing equation 7, we ignored order unity numerical coefficient (2.5 for our choice of and ). We discuss below in Section 3 how accounting for this coefficient would not change our final result. Planets were initially placed between 0.5 and 50 AU and we evolve their inward migration by integrating equation 6.
@@
where is set by equation 2, is a depletion factor we vary between 0.001 and 1, and Myr is the disk lifetime (e.g., E. E. Mamajek, 2009).444While the more improved disk lifetime is closer to 7–8 Myr (e.g., A. Michel et al., 2021), Jupiters are more likely to appear around more massive stars which tend to have shorter-lived disks (e.g., Á. Ribas et al., 2015) so we use 2 Myr as our nominal value.
@@ 2.2.1 Dust Replenishment
Our model setup implicitly treats the planet feeding zone to be isolated (i.e., no replenishment of solids). For planetesimal accretion, unless there are neighboring planets that can systematically scatter the planetesimals in, there is no a priori reason why a given feeding zone would be replenished. For pebble accretion, the picture is more complicated as the pebbles drift in aerodynamically and the drift timescale is shorter than pebble accretion time by factors of up to 10 (e.g., J. W. Lin et al., 2018; Y. Chachan & E. J. Lee, 2023). One way to stop this radial drift-in is to establish a pressure trap (e.g., P. Pinilla et al., 2011). While the exact source of the pressure trap remains debated—the most popular being perturbation by an embedded planet—it is a natural mechanism to explain both the longevity of dust disks and the observed concentric ring-like structures in protoplanetary disks (see, e.g., J. Bae et al., 2023, for a review).
@@
Given the range of possibilities, we explore the two limiting cases. Our isolated in situ pebble accretion corresponds to the case of a perfect trap created by the accreting Jupiter. We introduce another limiting case of a trap that is so inefficient that the trap is functionally non-existent (see e.g., B. Bitsch & J. Mah, 2023, for a more sophisticated disk model under this assumption). In this latter case, the planet has access to all the initial dust material exterior to its orbit:
@@ 2.2.2 Water Abundance
So far, we have computed the total mass of the metals that would be accreted onto a giant planet through post-formation pollution. Creating a large uniform sample of atmospheric measurements is most amenable for volatile species such as water which is more easily detectable than tracers of refractory species such as sulfur or nitrogen. Treatment of water abundance depends on the location at which the solids were accreted. Inside the iceline, water is in gaseous form, so post-formation accretion of solids would not impact water abundance, hence the water abundance is set by the initial solar value, which we take as 0.48 (oxygen mass fraction among all metals; N. Grevesse & A. Noels 1993) times 0.02 (our adopted solar metallicity), irrespective of pollution parameters. Outside the iceline, water is in solid form, so we estimate the water abundance by multiplying the total accreted metal mass by 0.48, assuming all the oxygen to be incorporated into water for simplicity (see, e.g., K. Heng et al., 2016, their Figure 3, “cold” case). We take the water iceline at 160 K which is 0.86 AU (291 days) based on our adopted disk temperature profile (equation 3). For more direct comparison with observable, we further multiply our mass ratio by 2.3/18 (solar metallicity gas mean molecular weight / mass of water molecule) to obtain number ratio. We focus solely on water abundance rather than C/O ratio because proper prediction of the latter requires careful accounting of carbon bearing species as well (e.g., C. Mordasini et al., 2016; E. Vlahos et al., 2024) which is beyond the scope of this paper.
@@ 3 Results
The predicted trends of metallicity (accreted solid mass over Jupiter mass) following post-formation pollution by planetesimal accretion, pebble accretion, and disk-induced migration are found in Figures 2, 3, and 4 respectively. The abundances shown in these figures represent what is added to the planet after formation; therefore, they do not represent the absolute atmospheric abundances—in order to obtain the absolute abundances, our results would have to be added to the metallicities the planets inherited at formation from the disk, which we assume to be stellar for simplicity. The implication is that any additive metallicity that falls below the stellar value (fixed to be 0.02 for simplicity) at a level that is smaller than what is detectable, would simply not show up in observations as a signature of pollution. Since both planetesimal accretion and migration produce additive metallicity that is an order of magnitude smaller than 0.02, we can reasonably rule them out as observable (even if we account for the order unity factor we ignored in the migration rate in case of pollution by migration). Instead, we focus on pebble accretion with high and high which produce sufficiently high additive metallicity. We have adopted a conservative case of planetesimal accretion here given the uncertainties introduced by the unknown dynamics of planetesimals—we revisit in Section 5 a case for a more optimistic scenario.
@@
which drops with orbital distance. In other words, the key difference is that in our conservative limit of planetesimal accretion the planet’s accretion cross section stays fixed, whereas in local pebble accretion the accretion cross section is bigger farther from the star.
Likewise, under disk-induced migration, metallicity declines with the final orbital period. By definition, if the final orbital period is long, there has been minimal radial migration and the planet spends most time where is lower, leading to overall lower metallicity by pollution. Figure 4 shows that the behavior of metallicity with respect to (gas disk depletion factor) is non-monotonic. While the total amount of metals accreted for planets migrating in gas-full () disks is larger, these planets also migrate the fastest and reach the shortest final orbital period, causing an overall leftward shift in the metallicity-final-period diagram (e.g., planets that initially start at 105 days with the lowest metallicity park at 103 days) with the planets that attain highest metallicity all piling up at the disk inner edge of 10 days.
@@ 4.1 Accuracy and Precision of Tier 2 Transit Spectroscopy Retrievals
To simulate Ariel observations using real targets, we use the July 2024 version of the Ariel Mission Candidate Sample (MCS) presented by B. Edwards & G. Tinetti (2022) and available on GitHub.555https://github.com/arielmission-space/Mission_Candidate_Sample We use the optimistic version of the MCS, which includes both confirmed and candidate planets. We conservatively adopt the easiest targets observable in a cumulative observing time of 1 year, where we assume that each transit observation takes 3 the transit duration and each target is re-observed until it reaches Tier 2 signal-to-noise. The number of observations required for each targets were pulled directly from the MCS. Our mock survey includes 431 Jupiters comprised of 172 confirmed and 259 candidate planets.
Ariel’s targets are hot and warm Jupiters around FGK and early M type dwarfs so they are all within the iceline. In the simplest baseline scenario of minimal migration after pollution, we expect the water abundance to be flat with respect to orbital period set to the stellar value irrespective of the existence/efficiency of a dust trap, as shown in Figure 5. We initially assign this stellar water abundance to each planet (adding stellar [Fe/H] to the solar value we adopt) and then randomly shift the abundance by 0.100.07 dex. Figure 6 depicts simulated Ariel retrievals of 431 Tier 2 transmission spectroscopy targets after 1 year of cumulative observation time with Ariel. A fit through all the planets (i.e., raw data without binning) recovers the expected flat H2O abundance with respect to orbital period.
@@
Under high-eccentricity migration, a gas giant first completes its formation at large orbital separations (beyond the iceline). After the disk dissipates, the giant is excited to a highly eccentric orbit by neighboring planetary or stellar perturbers, then tidally circularizes to a tight orbit with orbital period less than 10 days. Under all circumstances, giants that complete their pollution beyond the iceline are expected to have a significantly water-enriched envelope as compared to those that remain inside the iceline (because solid pollution adds extra water only beyond the iceline). We therefore expect the signature of high-eccentricity migration to show up as a sudden jump in water abundance ratio for hot Jupiters compared to warm Jupiters—insofar as most warm Jupiters did not undergo similar dynamical upheaval (e.g., M. Rice et al., 2022b; X.-Y. Wang et al., 2024).666For the purposes of the simulation, we assume that the migrating Jovian planets have already reached an orbital period on the order of days. Additionally, we assume that the metallicity of the planet remains constant throughout the high-eccentricity migration.
@@
Our goal is to assess the ability of Ariel survey at the Tier 2 level to discern the aforementioned effect of high-eccentricity migration on the measurable water abundance of hot Jupiters. To simulate the migration of a giant from beyond to inside the iceline, we assign the water abundance to the MCS targets with orbital periods inside 10 days to that drawn from an arbitrary distance between 1 and 10 AU, from the model expectation in Figure 5 corrected to non-solar abundance drawing from stellar [Fe/H] in the MCS, as well as uncertainties drawn from Q. Changeat & K. H. Yip (2023) set to 0.100.07 dex. The 10 AU upper limit to the orbital distance is motivated by the observed peak of gas giant occurrence rate at 1–10 AU (B. J. Fulton et al., 2021). The MCS targets with orbital periods longer than 10 days are assigned the model-expected water abundance at their measured periods (i.e., stellar value), accounting for the same stellar [Fe/H] correction and measurement uncertainties. Hot and warm Jupiters are separated, and each sorted to maximize the total number of targets observed. The final list is then a recombination of both, but with a fixed number ratio of 14 hot Jupiters for every 4 warm Jupiters, a ratio representing their relative occurence rate reported by E. A. Petigura et al. (2018).
Figure 7 illustrates the resulting water abundance of hot vs. warm Jupiters. As expected, polluted hot Jupiters that undergo high-eccentricity migration from beyond the iceline are expected to be significantly more water-enriched than warm Jupiters. The lower number of Tier 2 targets in this exercise compared to that shown in Figure 6 is a result of the additional constraint, wherein we now enforce a minimum number of warm Jupiters which predominantly require more observation time.
@@
Taking the weighted average between hot vs. warm Jupiters with the division at 10 days, the detection confidence exceeds 3–4. In case of no dust trap, the difference is visible by eye. The scenario with no dust traps predicts the hot Jupiters to be noticeably more water-rich than those with perfect dust traps simply because they have access to more solids (see Figure 5). Much of the scatter in true H2O value for hot Jupiters comes from our random distribution of the pre-migration orbital distance without any correlation to the final location in addition to the scatter in stellar metallicity. While a more detailed orbital evolution may beget further trends with e.g., orbital period among the hot Jupiters, we defer such studies to future investigation.